OpenStudy (jamesreed30):

Find the exact circumference in terms of PI of a circle with the given radius. R= 7 feet

OpenStudy (sfulks):

C = 2(pi)r

OpenStudy (jamesreed30):

It said that wasn't right

jhonyy9 (jhonyy9):

formula for circumference of a circle is 2*pi*r so than you substitute this value of r=7 in place of r inside this formula and multiply it by 2 so will get right sure the circumference of this circle with radius equal 7 feet in term of pi hope helped

OpenStudy (cal44):

The formula for the circumference of a circle is C=2(pi)r Since we know the radius r we can substitute C=2(pi)(7) since everything is being multiplied just switch it and simply to get C=14pi